A ball of mass 350g is kicked with a speed of 70 m/s. What is kinetic energy?
1
Expert's answer
2019-03-11T12:15:34-0400

The kinetic energy by definition


"K=\frac{mv^2}{2}"

"K=\frac{0.350\:\rm{kg}\times (70\:\rm{m/s})^2}{2}=857.5\:\rm{J}"